answer all questions

11 0 oints calculacians. Round your final answer to 2 decimal places.
Answer #1

Number of shares held = 100

Purchase price = $ 11 per share

Number of year shares held = 3 years

Sell Price = $ 17 per share

Annual Return = (Sale value / Purchase value) ^(1 +Number of year held) -1

= (17*100 / 11*100) ^(1/3) -1

= 0.1562 = 15.62%

OR

Simple return = (Sell price - Purchase price) /Purchase price = (17 -11) /11 = 54.55%

Annual Return = (Simple return +1) ^ (1 /Number of year held) -1

= (0.5455 +1) ^ (1/3) -1

= 0.1562 = 15.62%